# On the first-order genus of wreath products and their central extensions

### 16 Mar 2026 v1

@@ 1 Introduction
Theorem 1.1.
Let $A$ and $B$ be free abelian groups of finite ranks. Then $B\,{\rm wr}\,A$ is regularly injectively bi-interpretable with $\mathbb{Z}$.
@@
Let $A$ and $B$ be free abelian groups of finite ranks. Then $B\,{\rm wr}\,A$ is QFA.
@@
Recall that the space of finitely generated marked groups $\mathcal{G}$ can be informally defined as the set of all pairs $(G,A)$, where $G$ is a group and $A$ is an ordered finite generating set of $G$, endowed with the topology induced by local convergence of Cayley graphs. Given a finitely generated group $G$, we denote by $G]\subseteq\mathcal{G}$ its isomorphism class; that is,
@@
There are three immediate consequences of this result. The first one concerns the complexity of the isomorphism relation. Recall that an equivalence relation $E$ on a topological space $X$ is called smooth if there is a Polish space $P$ and a Borel map $\beta\colon X\to P$ such that for any $x,y\in X$, we have $xEy$ if and only if $\beta(x)=\beta(y)$.
In [46], Williams proved that the isomorphism relation on the space of $3$-step solvable, finitely generated, marked groups is non-smooth. We strengthen this result as follows.
@@ 2.1 Interpretability and bi-interpretability
In this case we denote $A$ by $\phi(B^{n})$ or $\phi({\mathbb{B}})$ and say that $\phi$ defines $A$ in $\mathbb{B}$. Similarly, an operation $f$ on the subset $A$ is definable in $\mathbb{B}$ if its graph is definable in $\mathbb{B}$. A constant $c$ is definable if the relation $x=c$ is definable. An $n$-ary predicate $P(x_{1},\ldots,x_{n})$ is definable in $\mathbb{B}$ if the set $\{(b_{1},\ldots,b_{n})\in\mathbb{B}^{n}|P(b_{1},\ldots,b_{n})\ {\rm is\ true}\}$ is definable in $\mathbb{B}$.
@@
In the notation above if $\sim$ is a definable equivalence relation on a definable subset $A\subseteq B^{n}$ then we say that the quotient set $A/\sim$ is interpretable in $\mathbb{B}$. Furthermore, an operation $f$ or a predicate $P$ on the quotient set $A/\sim$ is interpretable in $\mathbb{B}$ if the full preimage of its graph in $A$ is definable in $\mathbb{B}$. For example, if $N$ is a normal definable subgroup of a group $G$, then the equivalence relation $x\sim y$ on $G$ given by $xN=yN$ is definable in $G$, so the quotient set $G/N$ of all right cosets of $N$ is interpretable in $G$. It is easy to see that the multiplication induced from $G$ on $G/N$ is also interpretable in $G$. This shows that the quotient group $G/N$ is interpretable in $G$.
@@
From the arithmetical hierarchy and its relation to Turing degrees (usually referred to as Post’s theorem), it follows that every computably enumerable subset of ${\mathbb{N}}^{n}$, $n\geq 1$, is absolutely definable in $\mathbb{N}$} (see [41, 45]). This provides a very useful tool for interpretability of algebraic structures in $\mathbb{N}$}. To make it more precise, recall that an algebraic structure ${\mathcal{A}}=\langle A;L\rangle$ in a finite language $L$ is computable or recursive (see [39, 25]) if there is a surjective map $\nu:{\mathbb{N}}\to{\mathcal{A}}$, called an enumeration of $\mathcal{A}$}, such that the full preimages with respect to $\nu$ of all the graphs of the basic operations and predicates, including the equality =, on $\mathcal{A}$} are computable in $\mathbb{N}$}. For such an enumeration $\nu$ there is a computable function $\lambda:{\mathbb{N}}\to{\mathbb{N}}$, where $\lambda(i)=j$ if $j$ is the smallest natural number such that $\nu(j)=\nu(i)$. Put ${\mathbb{N}}_{0}=\lambda({\mathbb{N}})$. Then the restriction $\nu_{0}$ of $\nu$ to ${\mathbb{N}}_{0}$ gives a bijective enumeration $\nu_{0}:{\mathbb{N}}_{0}\to A$ such that the full $\nu_{0}$-preimages of the basic operations and predicates on $A$ are computable in $\mathbb{N}$}. Since $\nu_{0}$ is a bijection, this construction gives an $L$-structure denoted by ${\mathcal{A}}_{\nu}$ on ${\mathbb{N}}_{0}$ that is isomorphic to $\mathcal{A}$} with respect to $\nu_{0}$. Since the basic operations and predicates are computable in $\mathbb{N}$}, they are definable in $\mathbb{N}$}; hence, ${\mathcal{A}}_{\nu}$ is an interpretation of $\mathcal{A}$} in $\mathbb{N}$} with the coordinate map $\nu_{0}$.
@@ 2.3 Wreath products of free abelian groups
Let $A$ and $B$ be free abelian groups with bases $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$, respectively. If not stated otherwise, we view the groups $A$ and $B$ in multiplicative notation. Denote by $G=B\,{\rm wr}\,A$ the restricted wreath product of $A$ and $B$. We fix this notation throughout the entire paper.
By construction, the group $G$ is a semidirect product $N\rtimes A$, where $N$ is the direct product $N\simeq\Pi_{a\in A}B_{a}$ of isomorphic copies $B_{a}$ of $B$, termed the base group of this wreath product, and $A$, termed the active group, acts on $N$ by shifting the copies $B_{a}$ in $N$. Namely if $f\in\Pi_{a\in A}B_{a}$ is viewed as a function $f:A\to\cup_{a\in A}B_{a}$, then the action of $a\in A$ on $f$ results in a function $f^{a}(x)=f(ax)$ for any $x\in A$. As usual, elements $g\in G$ can be written uniquely in a normal form $g=af$, where $a\in A,f\in N$, and the multiplication is given by
@@
Note that in this case $f^{a}=a^{-1}fa$ - the conjugation of $f$ by $a$. The groups $A$ and $B$ canonically embed into $G$ via the maps $a\in A\to a\cdot 1_{N}\in G$ and $b\in B\to f_{b}$, where $f_{b}(x)=1$ for $x\neq 1$ and $f_{b}(1)=b$. We often identify $A$ and $B$ with their images in $G$, so $B$ is identified with the copy $B_{1}$.
@@
    The subgroup $N$ of $G$ is precisely the normal closure $N=ncl(B)$ of $B$ in $G$.
@@
    The action of $A$ on $N$ extends linearly to the action of the integer group ring ${\mathbb{Z}}A$ on $N$, so for $u=\Sigma n_{i}a_{i}\in{\mathbb{Z}}A$ and $t\in N$ $t^{u}=\Pi(t^{n_{i}})^{a_{i}}$. This makes $N$ into ${\mathbb{Z}}A$-module. Observe, that we use here the multiplicative notation for the action of ${\mathbb{Z}}A$ on $N$, which matches naturally with the conjugation notation in $G$.
@@
For convenience, in this and similar cases, we write sometimes $p=\bar{a}\cdot\bar{b}$, where $\bar{a}=(a_{1},\ldots,a_{m})$, $\bar{b}=(b_{1},\ldots,b_{n})$ and $\cdot$ is concatenation of tuples. Respectively, we write $s_{p}(g)=\bar{\gamma}\cdot\bar{P}$ for the tuple of coordinates of $g$ relative to $p$. Using this notation we sometimes write $g={\bar{a}}^{\bar{\gamma}}\cdot{\bar{b}}^{\bar{P}}$.
@@
    $C=\langle c_{1},\ldots,c_{m}\rangle\simeq A$, $G=C\cdot N$, $C\cap N=1$, and $N$ is a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$, when each $c_{i}$ acts on $N$ by conjugation.
@@
To verify that conditions 1), 2), and 3) above are equivalent, suppose first that an automorphism $\phi$ of $G$ maps $p$ to $q$. Then $\phi(A)=C$, so $C\simeq A$. By Lemma 2.7, item 2), $N$ is the centralizer of the commutant $G^{\prime}$ of $G$, hence $\phi(N)=N$. Clearly, $G=C\cdot N$ and $C\cap N=1$. The action of ${\mathbb{Z}}[a_{1},a_{1}^{-1},\ldots,a_{m},a_{m}^{-1}]$ on $N$ comes from the conjugation of the elements $a_{i}$ on $N$, so $\phi$ maps this action to the action of ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ on $N$, and $N$ becomes a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$, when each $c_{i}$ acts on $N$ by conjugation. Hence, 1) implies 2). Assume now that 2) holds; then $U=\langle u_{1},\ldots,u_{n}\rangle$ is isomorphic to $B$. Moreover, $G\simeq N\rtimes C$ and $G\simeq C\,{\rm wr}\,U$. Hence, the property 5) above holds for $G\simeq N\rtimes C$ and the formula (4) follows. This proves 3). To see that 3) implies 1) observe that if 3) holds then the coordinate-wise map $\psi_{0}:q\to p$ extends to a bijection on the normal forms
@@
Denote by $\Delta$ the augmentation ideal of ${\mathbb{Z}}A$, that is the kernel of the homomorphism ${\mathbb{Z}}A\to{\mathbb{Z}}$ induced by the trivial homomorphism $A\to 1$. As an ideal $\Delta$ is generated by elements $a_{1}-1,\ldots,a_{m}-1$ of ${\mathbb{Z}}A$. To see this note that $a_{i}^{-1}-1=\frac{1-a_{i}}{a_{i}}$ and uv1=u(v1)+(u1)uv-1=u(v-1)+(u-1) for $u,v\in A$. Put $\Delta^{0}=1$ and let $\Delta^{i}$ be the $i$th-power of $\Delta$ for $i>0$, i.e., it is an ideal in ${\mathbb{Z}}A$ generated by all products of the type $(y_{1}-1)\ldots(y_{i}-1)$, where $y_{j}\in\{a_{1},\ldots,a_{m}\}$. Denote by $N^{\Delta^{i}}$ the ${\mathbb{Z}}A$-submodule of $N$ generated by all elements $u^{P}$, where $u\in N$ and $P\in\Delta^{i}$. Put $R={\mathbb{Z}}A$, $R_{i}={\mathbb{Z}}A/\Delta^{i}$, and $N_{i}=N/N^{\Delta^{i}}$. In particular, $R_{0}\simeq R$, $R_{1}\simeq{\mathbb{Z}}$.
@@
To prove 1) recall that $G_{i}$ is generated as a normal subgroup by all simple left-normed commutators of the form $x_{1},\ldots,x_{i}]$ where $x_{k}\in\{a_{1},\ldots,a_{m},b_{1},\ldots,b_{n}\}$, $k=1,\ldots,i$ (this is true for any group $G$ and $x_{i}$ from any fixed generating set of $G$). Since $N$ is abelian and $G^{\prime}\leq N$ any commutator $x_{1},\ldots,x_{i}]$ where $x_{s}\in N$ for some $s>2$ is equal to 1. Therefore, $G_{i}$ is generated as a normal subgroup in $G$ by the commutators $b_{k},a_{j_{1}},\ldots,a_{j_{i-1}}]$, as claimed.
@@
For any non-trivial $g\in N$, every element in $N$ commutes with $g$ and every element in $A$ does not commute with $g$. Since elery element in $G$ is a product of an element in $N$ and an element in $A$, this gives 1) and 2). Since $A$ is abelian, for any $a\in A$, $A\subseteq C_{G}(a)$. And $a$ does not commute with any non-trivial $g\in N$, this gives 3).
@@ 3.2 Interpretability of $\mathbb{Z}$} in $G$
    The subgroup $N$ is 0-definable in $G$ by the formula
@@
    The subgroup $A$ is definable in $G$ with parameters by the formula
@@
    where $1\neq a$ is an arbitrary non-trivial element of $A$.
@@
Note first that in a free abelian group $A$ there is an automorphism $\theta\in\mathrm{Aut}A$ that maps $a$ to $a_{1}^{\beta}$ for some positive $\beta\in{\mathbb{N}}$. Replacing the basis $\{a_{1},\ldots,a_{m}\}$ of $A$ with basis $\{\theta(a_{1}),\dots,\theta(a_{m})\}$, if necessary, we may assume that $a=a_{1}^{\beta}$.
@@
Since $b_{1},\ldots,b_{n}$ is a basis of the free module $N$, one gets $Q_{i}=0$ for all $i>1$. Hence
@@
which again has the form (10). By induction, $M_{i}=M_{i+1}a_{1}^{\beta}$ for all $i=1,\ldots,k-1$. In particular, $g_{2}$ divides $g_{1}$. Since $g_{1}$ and $g_{2}$ do not have common divisors besides $1$ and 1-1, we deduce that $g_{2}=\pm 1$, and from $g=\frac{g_{1}}{g_{2}}$ we have $g_{2}=1$. It follows that $P_{1}=a_{1}^{\beta(k-1)}+\ldots a_{1}^{\beta}+1$ and $g=g_{1}=a^{k\beta}$ for some $k>0$. However, if we assume above that $g_{2}$ contains $a_{1}$. Then a similar argument shows that in this case $g_{1}=1$ and $g_{2}=a_{1}^{k\beta}$ for some $k>0$, so $g=a_{1}^{-k\beta}$, as claimed.
@@
Observe, that $G\models Base_{top}(a_{1},\ldots,a_{m})$, in fact, this it true for every basis $(a_{1}^{\prime},\ldots,a_{m}^{\prime})$ of $A$, because from the beginning $(a_{1},\ldots,a_{m})$ was chosen as an arbitrary basis of $A$. Suppose that $G\models Base_{top}(c_{1},\ldots,c_{m})$ for some elements $c_{1},\ldots,c_{m}\in G$. Then for every $c_{i}$ the formula $Cyc_{A}(c_{i},y)$ defines in $G$ an abelian subgroup $C_{i}$, which contains a non-trivial element $c_{i}$. These subgroups satisfy all the conditions 1)-6), hence
@@
From 7) it follows that $C$ is a free abelian of rank $m$, so every group $C_{i}$ is cyclic. Denote by $d_{i}u_{i}$, where $d_{i}\in A,u_{i}\in N$, a generator of $C_{i}$. It is easy to see that $d_{1},\dots,d_{m}$ form a basis of $A$. Therefore, there is an automorphism $\psi\in Aut(A)$ such that $\psi(a_{i})=d_{i},i=1,\ldots,m$. We claim that the automorphism $\psi$ extends to an automorphism $\hat{\psi}$ of $G$. Clearly, $\psi$ extends linearly to an automorphism $\lambda_{\psi}$ of ${\mathbb{Z}}A$. Now, $\lambda_{\psi}$ gives rise to a semilinear bijection $\tilde{\psi}$ of the ${\mathbb{Z}}A$-module $N$, where for $u=\Pi b_{i}^{P_{i}}\in N,P_{i}\in{\mathbb{Z}}A$,
@@
The isomorphism ${\tilde{\psi}}^{-1}$ maps $d_{i}u_{i}$ to $a_{i}v_{i}$, where $v_{i}={\tilde{\psi}}^{-1}(u_{i})\in N$, so $(a_{1}v_{1},\ldots,a_{m}v_{m})$ forms a basis of a free abelian group $A^{\prime}$, which is isopmorphic to $A$ via a map $\theta:a_{i}\to a_{i}v_{i},i=1,\ldots,n$. Note, that the elements $a_{i}v_{i}$ act by conjugation on $N$ precisely the same way as elements $a_{i}$, hence the map $au\to\theta(a)u$, for $a\in A,u\in N$ is an isomorphism of $G$. This shows that the tuples $(a_{1},\ldots,a_{m})$, $(a_{1},v_{1},\ldots,a_{m}v_{m})$, and $(d_{1}u_{1},\ldots,d_{m}u_{m})$ are automorphically equivalent in $G$.
Recall that $c_{i}\in C_{i}=\langle d_{i}u_{i}\rangle$, so $c_{i}=(d_{i}u_{i})^{\gamma_{i}}$, for some $\gamma_{i}\in{\mathbb{Z}}$. Now we apply Lemma 3.2 where the basis $a_{1},\ldots,a_{m}$, hence the group $A$, is replaced with the basis $(d_{1}u_{1},\ldots,d_{m}u_{m})$ and the group $A^{\prime}$, and the element $a$ is replaced by $c_{i}\in A^{\prime}$. This gives us that the formula $Cyc(c_{i},y)$ defines in $G$ the cyclic subgroup $\langle c_{i}\rangle$ which is equal to $C_{i}$. Therefore $c_{i}=(d_{i}u_{i})^{\pm 1}$, so there is an automorphism of $G$ that maps $c_{i}$ to $d_{i}u_{i}$ for $i=1,\ldots,m$. This implies that the tuples $(c_{1},\ldots,c_{m})$ and $(a_{1},\ldots,a_{m})$ are automorphically equivalent in $G$, as required.

### 18 Mar 2026 v2

@@ 1 Introduction
Theorem $1.$
Let AA and $B$ be free abelian groups of finite ranks. Then $B\,{\rm wr}\,A$ is regularly injectively bi-interpretable with $\mathbb{Z}$.
@@
Let AA and $B$ be free abelian groups of finite ranks. Then $B\,{\rm wr}\,A$ is QFA.
@@
Recall that the space of finitely generated marked groups $\mathcal{G}$ can be informally defined as the set of all pairs $(G,A)$, where $G$ is a group and AA is an ordered finite generating set of $G$, endowed with the topology induced by local convergence of Cayley graphs. Given a finitely generated group $G$, we denote by $G]\subseteq\mathcal{G}$ its isomorphism class; that is,
@@
There are three immediate consequences of this result. The first one concerns the complexity of the isomorphism relation. Recall that an equivalence relation $E$ on a topological space $X$ is called smooth if there is a Polish space $P$ and a Borel map $\beta\colon X\to P$ such that for any $x,y\in X$, we have xEyxEy if and only if $\beta(x)=\beta(y)$.
In [46], Williams proved that the isomorphism relation on the space of 33-step solvable, finitely generated, marked groups is non-smooth. We strengthen this result as follows.
@@ 2.1 Interpretability and bi-interpretability
In this case we denote AA by $\phi(B^{n})$ or $\phi({\mathbb{B}})$ and say that $\phi$ defines AA in $\mathbb{B}$. Similarly, an operation $f$ on the subset AA is definable in $\mathbb{B}$ if its graph is definable in $\mathbb{B}$. A constant $c$ is definable if the relation $x=c$ is definable. An $n$-ary predicate $P(x_{1},\ldots,x_{n})$ is definable in $\mathbb{B}$ if the set $\{(b_{1},\ldots,b_{n})\in\mathbb{B}^{n}|P(b_{1},\ldots,b_{n})\ {\rm is\ true}\}$ is definable in $\mathbb{B}$.
@@
In the notation above if $\sim$ is a definable equivalence relation on a definable subset $A\subseteq B^{n}$ then we say that the quotient set $A/\sim$ is interpretable in $\mathbb{B}$. Furthermore, an operation $f$ or a predicate $P$ on the quotient set $A/\sim$ is interpretable in $\mathbb{B}$ if the full preimage of its graph in AA is definable in $\mathbb{B}$. For example, if NN is a normal definable subgroup of a group $G$, then the equivalence relation $x\sim y$ on $G$ given by $xN=yN$ is definable in $G$, so the quotient set $G/N$ of all right cosets of NN is interpretable in $G$. It is easy to see that the multiplication induced from $G$ on $G/N$ is also interpretable in $G$. This shows that the quotient group $G/N$ is interpretable in $G$.
@@
From the arithmetical hierarchy and its relation to Turing degrees (usually referred to as Post’s theorem), it follows that every computably enumerable subset of ${\mathbb{N}}^{n}$, $n\geq 1$, is absolutely definable in $\mathbb{N}$} (see [41, 45]). This provides a very useful tool for interpretability of algebraic structures in $\mathbb{N}$}. To make it more precise, recall that an algebraic structure ${\mathcal{A}}=\langle A;L\rangle$ in a finite language $L$ is computable or recursive (see [39, 25]) if there is a surjective map $\nu:{\mathbb{N}}\to{\mathcal{A}}$, called an enumeration of $\mathcal{A}$}, such that the full preimages with respect to $\nu$ of all the graphs of the basic operations and predicates, including the equality =, on $\mathcal{A}$} are computable in $\mathbb{N}$}. For such an enumeration $\nu$ there is a computable function $\lambda:{\mathbb{N}}\to{\mathbb{N}}$, where $\lambda(i)=j$ if $j$ is the smallest natural number such that $\nu(j)=\nu(i)$. Put ${\mathbb{N}}_{0}=\lambda({\mathbb{N}})$. Then the restriction $\nu_{0}$ of $\nu$ to ${\mathbb{N}}_{0}$ gives a bijective enumeration $\nu_{0}:{\mathbb{N}}_{0}\to A$ such that the full $\nu_{0}$-preimages of the basic operations and predicates on AA are computable in $\mathbb{N}$}. Since $\nu_{0}$ is a bijection, this construction gives an $L$-structure denoted by ${\mathcal{A}}_{\nu}$ on ${\mathbb{N}}_{0}$ that is isomorphic to $\mathcal{A}$} with respect to $\nu_{0}$. Since the basic operations and predicates are computable in $\mathbb{N}$}, they are definable in $\mathbb{N}$}; hence, ${\mathcal{A}}_{\nu}$ is an interpretation of $\mathcal{A}$} in $\mathbb{N}$} with the coordinate map $\nu_{0}$.
@@ 2.3 Wreath products of free abelian groups
Let AA and $B$ be free abelian groups with bases $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$, respectively. If not stated otherwise, we view the groups AA and $B$ in multiplicative notation. Denote by $G=B\,{\rm wr}\,A$ the restricted wreath product of AA and $B$. We fix this notation throughout the entire paper.
By construction, the group $G$ is a semidirect product $N\rtimes A$, where NN is the direct product $N\simeq\Pi_{a\in A}B_{a}$ of isomorphic copies $B_{a}$ of $B$, termed the base group of this wreath product, and AA, termed the active group, acts on NN by shifting the copies $B_{a}$ in NN. Namely if $f\in\Pi_{a\in A}B_{a}$ is viewed as a function $f:A\to\cup_{a\in A}B_{a}$, then the action of $a\in A$ on $f$ results in a function $f^{a}(x)=f(ax)$ for any $x\in A$. As usual, elements $g\in G$ can be written uniquely in a normal form $g=af$, where $a\in A,f\in N$, and the multiplication is given by
@@
Note that in this case $f^{a}=a^{-1}fa$ - the conjugation of $f$ by $a$. The groups AA and $B$ canonically embed into $G$ via the maps $a\in A\to a\cdot 1_{N}\in G$ and $b\in B\to f_{b}$, where $f_{b}(x)=1$ for $x\neq 1$ and $f_{b}(1)=b$. We often identify AA and $B$ with their images in $G$, so $B$ is identified with the copy $B_{1}$.
@@
    The subgroup NN of $G$ is precisely the normal closure $N=ncl(B)$ of $B$ in $G$.
@@
    The action of AA on NN extends linearly to the action of the integer group ring ${\mathbb{Z}}A$ on NN, so for $u=\Sigma n_{i}a_{i}\in{\mathbb{Z}}A$ and $t\in N$ $t^{u}=\Pi(t^{n_{i}})^{a_{i}}$. This makes NN into ${\mathbb{Z}}A$-module. Observe, that we use here the multiplicative notation for the action of ${\mathbb{Z}}A$ on NN, which matches naturally with the conjugation notation in $G$.
@@
For convenience, in this and similar cases, we write sometimes $p=\bar{a}\cdot\bar{b}$, where $\bar{a}=(a_{1},\ldots,a_{m})$, $\bar{b}=(b_{1},\ldots,b_{n})$ and $\cdot$ is concatenation of tuples. Respectively, we write $s_{p}(g)=\bar{\gamma}\cdot\bar{P}$ for the tuple of coordinates of $g$ relative to pp. Using this notation we sometimes write $g={\bar{a}}^{\bar{\gamma}}\cdot{\bar{b}}^{\bar{P}}$.
@@
    $C=\langle c_{1},\ldots,c_{m}\rangle\simeq A$, $G=C\cdot N$, $C\cap N=1$, and NN is a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$, when each $c_{i}$ acts on NN by conjugation.
@@
To verify that conditions 1), 2), and 3) above are equivalent, suppose first that an automorphism $\phi$ of $G$ maps $p$ to $q$. Then $\phi(A)=C$, so $C\simeq A$. By Lemma 2.7, item 2), NN is the centralizer of the commutant $G^{\prime}$ of $G$, hence $\phi(N)=N$. Clearly, $G=C\cdot N$ and $C\cap N=1$. The action of ${\mathbb{Z}}[a_{1},a_{1}^{-1},\ldots,a_{m},a_{m}^{-1}]$ on NN comes from the conjugation of the elements $a_{i}$ on NN, so $\phi$ maps this action to the action of ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ on NN, and NN becomes a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$, when each $c_{i}$ acts on NN by conjugation. Hence, 1) implies 2). Assume now that 2) holds; then $U=\langle u_{1},\ldots,u_{n}\rangle$ is isomorphic to $B$. Moreover, $G\simeq N\rtimes C$ and $G\simeq C\,{\rm wr}\,U$. Hence, the property 5) above holds for $G\simeq N\rtimes C$ and the formula (4) follows. This proves 3). To see that 3) implies 1) observe that if 3) holds then the coordinate-wise map $\psi_{0}:q\to p$ extends to a bijection on the normal forms
@@
Denote by $\Delta$ the augmentation ideal of ${\mathbb{Z}}A$, that is the kernel of the homomorphism ${\mathbb{Z}}A\to{\mathbb{Z}}$ induced by the trivial homomorphism $A\to 1$. As an ideal $\Delta$ is generated by elements $a_{1}-1,\ldots,a_{m}-1$ of ${\mathbb{Z}}A$. To see this note that $a_{i}^{-1}-1=\frac{1-a_{i}}{a_{i}}$ and uv1=u(v1)+(u1)uv-1=u(v-1)+(u-1) for $u,v\in A$. Put $\Delta^{0}=1$ and let $\Delta^{i}$ be the $i$th-power of $\Delta$ for $i>0$, i.e., it is an ideal in ${\mathbb{Z}}A$ generated by all products of the type $(y_{1}-1)\ldots(y_{i}-1)$, where $y_{j}\in\{a_{1},\ldots,a_{m}\}$. Denote by $N^{\Delta^{i}}$ the ${\mathbb{Z}}A$-submodule of NN generated by all elements $u^{P}$, where $u\in N$ and $P\in\Delta^{i}$. Put $R={\mathbb{Z}}A$, $R_{i}={\mathbb{Z}}A/\Delta^{i}$, and $N_{i}=N/N^{\Delta^{i}}$. In particular, $R_{0}\simeq R$, $R_{1}\simeq{\mathbb{Z}}$.
@@
To prove 1) recall that $G_{i}$ is generated as a normal subgroup by all simple left-normed commutators of the form $x_{1},\ldots,x_{i}]$ where $x_{k}\in\{a_{1},\ldots,a_{m},b_{1},\ldots,b_{n}\}$, $k=1,\ldots,i$ (this is true for any group $G$ and $x_{i}$ from any fixed generating set of $G$). Since NN is abelian and $G^{\prime}\leq N$ any commutator $x_{1},\ldots,x_{i}]$ where $x_{s}\in N$ for some $s>2$ is equal to 1. Therefore, $G_{i}$ is generated as a normal subgroup in $G$ by the commutators $b_{k},a_{j_{1}},\ldots,a_{j_{i-1}}]$, as claimed.
@@
For any non-trivial $g\in N$, every element in NN commutes with $g$ and every element in AA does not commute with $g$. Since elery element in $G$ is a product of an element in NN and an element in AA, this gives 1) and 2). Since AA is abelian, for any $a\in A$, $A\subseteq C_{G}(a)$. And $a$ does not commute with any non-trivial $g\in N$, this gives 3).
@@ 3.2 Interpretability of $\mathbb{Z}$} in $G$
    The subgroup NN is 0-definable in $G$ by the formula
@@
    The subgroup AA is definable in $G$ with parameters by the formula
@@
    where $1\neq a$ is an arbitrary non-trivial element of AA.
@@
Note first that in a free abelian group AA there is an automorphism $\theta\in\mathrm{Aut}A$ that maps $a$ to $a_{1}^{\beta}$ for some positive $\beta\in{\mathbb{N}}$. Replacing the basis $\{a_{1},\ldots,a_{m}\}$ of AA with basis $\{\theta(a_{1}),\dots,\theta(a_{m})\}$, if necessary, we may assume that $a=a_{1}^{\beta}$.
@@
Since $b_{1},\ldots,b_{n}$ is a basis of the free module NN, one gets $Q_{i}=0$ for all $i>1$. Hence
@@
which again has the form (10). By induction, $M_{i}=M_{i+1}a_{1}^{\beta}$ for all $i=1,\ldots,k-1$. In particular, $g_{2}$ divides $g_{1}$. Since $g_{1}$ and $g_{2}$ do not have common divisors besides 11 and 1-1, we deduce that $g_{2}=\pm 1$, and from $g=\frac{g_{1}}{g_{2}}$ we have $g_{2}=1$. It follows that $P_{1}=a_{1}^{\beta(k-1)}+\ldots a_{1}^{\beta}+1$ and $g=g_{1}=a^{k\beta}$ for some $k>0$. However, if we assume above that $g_{2}$ contains $a_{1}$. Then a similar argument shows that in this case $g_{1}=1$ and $g_{2}=a_{1}^{k\beta}$ for some $k>0$, so $g=a_{1}^{-k\beta}$, as claimed.
@@
Observe, that $G\models Base_{top}(a_{1},\ldots,a_{m})$, in fact, this it true for every basis $(a_{1}^{\prime},\ldots,a_{m}^{\prime})$ of AA, because from the beginning $(a_{1},\ldots,a_{m})$ was chosen as an arbitrary basis of AA. Suppose that $G\models Base_{top}(c_{1},\ldots,c_{m})$ for some elements $c_{1},\ldots,c_{m}\in G$. Then for every $c_{i}$ the formula $Cyc_{A}(c_{i},y)$ defines in $G$ an abelian subgroup $C_{i}$, which contains a non-trivial element $c_{i}$. These subgroups satisfy all the conditions 1)-6), hence
@@
From 7) it follows that $C$ is a free abelian of rank $m$, so every group $C_{i}$ is cyclic. Denote by $d_{i}u_{i}$, where $d_{i}\in A,u_{i}\in N$, a generator of $C_{i}$. It is easy to see that $d_{1},\dots,d_{m}$ form a basis of AA. Therefore, there is an automorphism $\psi\in Aut(A)$ such that $\psi(a_{i})=d_{i},i=1,\ldots,m$. We claim that the automorphism $\psi$ extends to an automorphism $\hat{\psi}$ of $G$. Clearly, $\psi$ extends linearly to an automorphism $\lambda_{\psi}$ of ${\mathbb{Z}}A$. Now, $\lambda_{\psi}$ gives rise to a semilinear bijection $\tilde{\psi}$ of the ${\mathbb{Z}}A$-module NN, where for $u=\Pi b_{i}^{P_{i}}\in N,P_{i}\in{\mathbb{Z}}A$,
@@
The isomorphism ${\tilde{\psi}}^{-1}$ maps $d_{i}u_{i}$ to $a_{i}v_{i}$, where $v_{i}={\tilde{\psi}}^{-1}(u_{i})\in N$, so $(a_{1}v_{1},\ldots,a_{m}v_{m})$ forms a basis of a free abelian group $A^{\prime}$, which is isopmorphic to AA via a map $\theta:a_{i}\to a_{i}v_{i},i=1,\ldots,n$. Note, that the elements $a_{i}v_{i}$ act by conjugation on NN precisely the same way as elements $a_{i}$, hence the map $au\to\theta(a)u$, for $a\in A,u\in N$ is an isomorphism of $G$. This shows that the tuples $(a_{1},\ldots,a_{m})$, $(a_{1},v_{1},\ldots,a_{m}v_{m})$, and $(d_{1}u_{1},\ldots,d_{m}u_{m})$ are automorphically equivalent in $G$.
Recall that $c_{i}\in C_{i}=\langle d_{i}u_{i}\rangle$, so $c_{i}=(d_{i}u_{i})^{\gamma_{i}}$, for some $\gamma_{i}\in{\mathbb{Z}}$. Now we apply Lemma 3.2 where the basis $a_{1},\ldots,a_{m}$, hence the group AA, is replaced with the basis $(d_{1}u_{1},\ldots,d_{m}u_{m})$ and the group $A^{\prime}$, and the element $a$ is replaced by $c_{i}\in A^{\prime}$. This gives us that the formula $Cyc(c_{i},y)$ defines in $G$ the cyclic subgroup $\langle c_{i}\rangle$ which is equal to $C_{i}$. Therefore $c_{i}=(d_{i}u_{i})^{\pm 1}$, so there is an automorphism of $G$ that maps $c_{i}$ to $d_{i}u_{i}$ for $i=1,\ldots,m$. This implies that the tuples $(c_{1},\ldots,c_{m})$ and $(a_{1},\ldots,a_{m})$ are automorphically equivalent in $G$, as required.